OpenStudy (mariamikayla):

1+3+3^2+....+3^49 is 3^49 -1 or 3^50 -1 ?

OpenStudy (mariamikayla):

\[3^{49}-1 \ \ or \ \ 3^{50} -1\]

OpenStudy (agent0smith):

1+3+3^2+....+3^49 already includes 3^49, plus 3^48 and a bunch of other terms far greater than 1... so how could it possibly equal 3^49 - 1?

OpenStudy (mariamikayla):

Using geometrical progression it gives me http://prntscr.com/bsgm2y it is right?

OpenStudy (agent0smith):

I think you might want to post a screenshot of the question.

OpenStudy (agent0smith):

1+3+3^2+....+3^49 Is a geometric series with a1 = 1 and r = 3 and n=50. Your original question makes very little sense due to your wording. If you need the sum of 1+3+3^2+....+3^49, then your work http://prntscr.com/bsgm2y looks fine

OpenStudy (mariamikayla):

Thanks a lot :D i've got it now ;)

OpenStudy (sshayer):

\[1+3+3^2+...+3^{49}=3^0+3^1+3^2+...+3^{49}\] number of terms=50 \[a=1,r=3,s _{50}=a \left( \frac{ r^n-1 }{ r-1 } \right)=1 \left( \frac{ 3^{50}-1 }{ 3-1 } \right)=?\]